Roz is twice as old as Grace. In 5 years the sum of their ages will be 28. How old are they now?
Found 2 solutions by maribeth_abella, Earlsdon:
Answer by maribeth_abella(11)

x = Grace age
2x = Roz age
x+5+2x+5 = 28
3x + 10 = 28
3x = 28 - 10
x = 28/3
x = 6

x = 6
2(6) = 12

Roz is 12 years old
Grace is 6 years old

Answer by Earlsdon(6294)

Let R be Roz's present age and G be Grace's present age.
1) R = 2G
2) R+5 + G+5 = 28 Substitute R = 2G and solve for G.
2G+5 + G+5 = 28 Simplify and solve for G.
3G + 10 = 28 Subtract 10 from both sides.
3G = 18 Divide both sides by 3.
G = 6
R = 2G
R = 12
Roz is 12 years old.
Grace is 6 years old.